the length of a rectangle is two less than twice its width.If the perimeter of the rectangle is 44 sq. units what are the dimensions of the rectangle?









Answer by mananth(16946)

width = w
length = 2w-2
..
perimeter = 44
..
2*(l+w)=perimeter
2*(w+2w-2)=44
2*(3w-2)=44
6w-4=44
+4 +4
6w=48
/6
w=8 units.
...
Length = 2w-2
2*8-2
=14 units
...
m.ananth@hotmail.ca